In Millikan’s oil-drop experiment, one looks at a small oil drop held motionless between two plates. Take the voltage between the plates to be 2033 V, and the plate separation to be 2.00 cm. The oil drop (of density 0.81 g/cm3) has a diameter of 4.7 x 10-6m. Find the charge on the drop, in terms of electron units. You need to round your answer to the nearest integer.

Oil droplet
Answer #1

Famg, F= 2€, E=Vd F = 98 = 24 mg = qu [d=2cm= 2x10 m g=9.8m/s2 q- med V=20334 mass m = (Bersih wolume) Density = 0.818/6m3 =